Prove that a subset of a set of measure zero has measure zero.


1
Expert's answer
2022-03-14T17:48:31-0400

By the properties of the measure "\mu" :

"A\subset B\Rightarrow \mu A\leq\mu B"

But "\mu C\geq0" for any measurable set "C". So, if the measure of the set "B" is zero:

"0\leq\mu A\leq\mu B=0\Rightarrow\mu A=0"